OpenStudy (anonymous):

Find the absolute maximum and absolute minimum values of the function f(x)=x^4−10x^2+3 on each of the indicated intervals. Enter -1000 for any absolute extrema that does not exist. (A) Interval = [−3,−1]. Absolute maximum = Absolute minimum = (B) Interval = [−4,1]. Absolute maximum = Absolute minimum = (C) Interval = [−3,4]. Absolute maximum = Absolute minimum

OpenStudy (campbell_st):

so have you got the 1st derivative and solved to the stationary points...?

OpenStudy (anonymous):

what do you mean by that

OpenStudy (campbell_st):

ok... well its hard to identify max and min values if you can't differentiate... stationary points are when a tangent has a zero gradient... sometimes called critical values...

OpenStudy (anonymous):

yeah i already have my critical points they are sqrt5,-sqrt5 and 0

OpenStudy (campbell_st):

ok... so you need to substitute x = -1, x = -sqrt(5) and x = -3 into the original equation to see which has the largest y value that is how you find the answer to part (a)

OpenStudy (campbell_st):

an easy way to do this problem is to use some graphing software... means no calculus...

OpenStudy (anonymous):

im still a bit confuse do i plug in these values into the derivtive and also where did these values come from

OpenStudy (campbell_st):

no plug then into the original equation... y = (-3)^4 - 10(-3)^2 + 3 = 81 - 90 + 3 y = -6 x = - sqrt(5) y = (-sqrt(5)^4 - 10(-sqrt(5)^2 + 3 y = 25 - 50 + 3 y = -22 x = -1 so the absolute maximum are x = -3 and x = -1 when y = -6 y = (-1)^4 - 10(-1)^2 + 3 y = 1 - 10 + 3 y = -6

OpenStudy (campbell_st):

and absolute min is y = -23 when x = -sqrt{5}

OpenStudy (anonymous):

does it matter in what interval i put it

OpenStudy (anonymous):

i put that as the answer for part a as the minimum but it says its wrong

OpenStudy (campbell_st):

where here is a graph of your function

OpenStudy (anonymous):

thank you hopefully i get t right lol

OpenStudy (campbell_st):

well its interesting it doesn't see y = -22 as the min on the interval [-3, -1]

OpenStudy (campbell_st):

hope it helps...

OpenStudy (anonymous):

dude ask her for a nudes unless shes like 10 years old shes a dumb bimbo

OpenStudy (anonymous):

ill try to ignore that comment but thanks again :) campell

OpenStudy (anonymous):

bye bimbo face